Use properties of Boolean algebra to simplify the following Boolean expression (showing all the steps):


(y′)(x′ + y′)′



1
Expert's answer
2021-11-12T03:56:07-0500

Use properties of Boolean algebra let us simplify the following Boolean expression:


"(x + y')(x' + y')'=
(x + y')(x'' \cdot y'')=
(x + y')(x\cdot y)\\=
x\cdot x\cdot y + y'\cdot x\cdot y=
x\cdot y + y'\cdot y\cdot x=
x\cdot y +0\cdot x\\=
x\cdot y +0=x\cdot y"